PREX/CREX Monday, Oct 17, 2016 3:00pm EDT

From PREX Wiki
Jump to: navigation, search

Back to Main Page >> PREX/CREX_Teleconferences

previous meeting << >> following meeting

Logistic information

 BlueJeans calling instructions:
 Toll-Free Number (U.S.&  Canada):  888-240-2560
 International toll number:         408-740-7256
 PARTICIPANT CODE:                  345647893#
 JLab meeting room: "CC L207"

Agenda

  1. Monitors- first look - Caryn
    • CP: we took several solid target runs @ 2.2 GeV
    • CP: different HV settings for SAMs (some of them good)
    • CP: we can check linearity of SAMs with large charge asymmetry, maybe calibrating BPM pedestals
    • CP: Dustin swapped PMTs and bases but the SAM signals remain large as in the spring (SAM6 is larger than SAM2). So it’s either reflectivities in the signals or more scattering on one side. SAM4 now is bigger than SAM8 swapping magnitudes from the spring.
    • KK: why do we worry about the level the signal? CP: maybe relevant to saturation DM: these are more curiosities but are not big problems, not sure how to properly explain it. KK: if you are off by 0.5 cm it could just be rate.
    • CP: cavities seem to show signal needs to be investigate CG: correlate that to the BPMs
    • CP: BPMs seemed to be fine with the twiddle scans but they look bad now (see att 11,12 haplog 3309) KP: we need a report on this KP: Caryn should analyzing and report on this
  2. BCM calibration - Caryn
    • CP: Trent reconfigured two of the outputs with more filters because they had some problems in the Gmp analysis
    • CP: however after MCC fixed their phase matching in the machine the unfiltered signal also does not have the DC level shift
    • CG: we should ask Mark for the latest results from his studies with the unfiltered digital receiver
    • CP: the BCMs are calibrated KP: take out the lower points and fit the higher currents because the analog seems to show a hump CP: will do
  3. PITA scan - Caryn
    • CP: PITA slope was 2.8 ppm per unit change last time, now it’s 0.028 ppm per unit change (100x too small)
    • KP: one obvious thing would if we made the correct changes.
    • CP: rotate the HWP by 22.5 deg and see what is happening? KP: sure
    • KK: get Joe to talk about this. CG: will do
  4. target chamber design - Kent :
    • meeting on phone on Friday with Silviu, Bobm Dave and Al
    • Silviu showed some pictures of the new design for the movers
    • Dave is happier with the Wayne/Silviu design
    • UVa cad was helpful; things are moving ahead
  5. optics simulation - Tyler:
    • TK: sent Kent some diagonistic plots KP: new envelopes TK: 5 degree track, also sent updated fits to the envelope to Nickies rootfiles (which have the correct collimator cuts)
    • TK: removed a bug from his simulation should be able to get any envelope we want
    • KP: these new envelopes look much better (will see what Chris/Adam say) .. will have to look at Q1
    • KP: you will be able to match his old FOM TK: will do
    • KK the collimator cuts fix our problem and we should get PREX vs energy FOM plots .. TK: after implement fix
    • KP: need 2d hit patters at different z TK: will do
  6. Student help with data taking - Julliete:
    • Sakihb is coming over for shift (Gmp)
    • CP: will be there Sat day and meeting him Friday evening to show him

Attendance

Ciprian G, Sanghwa P, Tao Y, Kent P, Caryn P, Paul S, Rakitha B, Seamus R, Juliette, KK, Dustin, Guido,Tyler,

Excused: